a boat travels 2 km/hr in still water. the boat travels 4 km in the same amount of time it takes to travel 1 km upstream. what is the speed of the river?
Answer by mananth(16949)

a boat travels 2 km/hr in still water. the boat travels 4 km in the same amount of time it takes to travel 1 km upstream. what is the speed of the river?
..
speed of boat in still water = 2km/h
..
let speed of current = c km/h
..
upstream seed 2-c
downstream speed = 2+c
..
upstream d=1
downstream d=4
..
t=d/r
t upstream = t downstream
1/(2-c)=4/(2+c)
cross multiply
(2+c)=4(2-c)
2+c=8-4c
5c=6
c=6/5 km/h
...
m.ananth@hotmail.ca